# Revisiting Locally Differentially Private Protocols: Towards Better Trade-offs in Privacy, Utility, and Attack Resistance

### v2

@@ V-D1 Adaptive Subset Selection (ASS)
$&\underset{\omega\in\mathbb{Z}}{\mathrm{min}}&&J(\omega)=w_{\mathrm{MSE}}\cdot\Bigg{(}\frac{(k-1)(k+2\omega e^{\varepsilon}-\omega)}{n\omega\left(-k+\omega+(k-1)e^{\varepsilon}-(\omega-1)e^{\varepsilon}\right)^{2}}\\ &&&\left.+\frac{\left(k-\omega+(\omega-1)e^{\varepsilon}\right)\left(-\omega(k-\omega)-\omega(\omega-1)e^{\varepsilon}\right)}{n\omega\left(-k+\omega+(k-1)e^{\varepsilon}-(\omega-1)e^{\varepsilon}\right)^{2}}\right)\\ &&&+w_{\text{ASR}}\cdot\left(\frac{e^{\varepsilon}}{\omega e^{\varepsilon}+k-\omega}\right),\\ &\emph{s.t.}&&1\leq\omega<k,\quad w_{\text{ASR}}+w_{\mathrm{MSE}}=1\mathrm{.}\end{aligned}$
@@
V-D2 Adaptive Unary Encoding (AUE)
@@
$+w_{\text{ASR}}\cdot\Bigg{(}(1-p)(1-q)^{k-1}\cdot\frac{1}{k}$
$\quad+\sum_{m=1}^{k}p\cdot\frac{1}{m}\cdot\binom{k-1}{m-1}q^{m-1}(1-q)^{(k-1)-(m-1)}\Bigg{)}\mathrm{,}$
s.t. $0.5\leq p<1,\;q=\frac{p}{e^{\varepsilon}(1-p)+p},\;w_{\text{ASR}}+w_{\mathrm{MSE}}=1\mathrm{.}$
@@ V-D3 Adaptive LH (ALH)
$\quad+w_{\text{ASR}}\cdot\left(\frac{e^{\varepsilon}}{(e^{\varepsilon}+g-1)\cdot\max\left\{\frac{k}{g},1\right\}}\right)\mathrm{,}$
@@
V-D4 Adaptive THE (ATHE)
The ATHE mechanism extends the THE protocol. Unlike THE [22], which only aims to minimize the MSE in Equation (20), ATHE jointly optimizes both MSE and ASR by adapting the threshold parameter $\theta$. Formally, the optimization problem for ATHE is defined as:
@@ V-D4 Adaptive THE (ATHE)
s.t. $0.5\leq\theta\leq 1,\;w_{\text{ASR}}+w_{\mathrm{MSE}}=1\mathrm{,}$
in which $p$ and $q$ are given in Equation (18). The constraint $0.5\leq\theta\leq 1$ follows the settings used in prior work [22].
@@ VI Experiments and Analysis
The objective of our experiments is to thoroughly evaluate the performance of the proposed adaptive LDP protocols in terms of privacy, utility and resilience against data reconstruction attacks. Specifically, in Section VI-B, we conduct an ASR analysis for each LDP protocol to quantify their vulnerability to data reconstruction attacks, thereby assessing the privacy guarantees offered by existing and newly proposed mechanisms. Subsequently, in Section VI-C, we perform an MSE analysis to evaluate the utility guarantees provided by existing and our newly proposed mechanisms. Next, in Section VI-D, we explore the trade-off between ASR and MSE (i.e., Pareto frontier), highlighting how our adaptive protocols compare to traditional ones in balancing privacy and utility under different scenarios. We then assess in Section VI-E the impact of different weights $w_{\text{ASR}}$ and $w_{\mathrm{MSE}}$ in the objective function to provide insights into the influence of prioritizing privacy versus utility. Afterwards, we briefly analyze in Section VI-F the parameter optimization approach for each adaptive protocol.
We highlight that to systematize these aforementioned experiments, we rely primarily on the analytical closed-form equations for both MSE and ASR, as prior research has demonstrated a strong agreement between analytical and empirical results [22, 28, 12]. This choice allows us to comprehensively evaluate the performance of our protocols across a wide range of scenarios while avoiding computational overhead. Nonetheless, we also include a concise empirical validation in Section VI-G and Appendix -D to ensure consistency and reinforce the validity of our analytical findings.
@@ VI-A Setup of Analytical Experiments
  • Environment. All algorithms are implemented in Python 3 and run on a desktop machine with 3.2GHz Intel Core i9 and 64GB RAM. All our code will be open-sourced.
  • LDP protocols. We experiment with the eight LDP protocols described in Section IV and our four adaptive LDP protocols described in Section V-D.
  • Number of users. For the analytical variance/MSE derived from Equation (4) (e.g., Equation (6) for GRR, Equation (20) for THE, …), we report variance per user, corresponding to the analytical derivation of $\text{Var}[]/n$. This allows us to examine the fundamental properties of each protocol independently of the dataset size.
@@
      High privacy regime with $\varepsilon\in\{0.5,0.6,\ldots,1.9,2.0\}$.
@@
  • Weights for optimization. Unless otherwise mentioned, we fix the weights for the two-objective optimization to $w_{\text{ASR}}=w_{\mathrm{MSE}}=0.5$, aiming for a balanced trade-off between privacy and accuracy. Experiments in Section VI-E will focus on varying these weights.
  • Optimization method. Parameters for our adaptive protocols (e.g., subset size $\omega$ for ASS) were tuned via grid search [43], followed by selection using a Weighted Scalarization strategy (see Section V-C). While grid search is exhaustive and ensures broad coverage of the parameter space, it can be computationally expensive. For large or high-dimensional search spaces, we recommend constrained or gradient-free optimization methods [45] as a more efficient alternative to accelerate convergence without significantly sacrificing solution quality.
@@ VI-B ASR Analysis for LDP Protocols
Figure 2: Attacker Success Rate (ASR) vs. privacy budget ($\varepsilon$) for different LDP frequency estimation protocols across varying domain sizes ($k$). The plots compare state-of-the-art LDP protocols, including GRR, SUE, BLH, OUE, OLH, SS, SHE and THE, against our newly proposed adaptive protocols (ASS, AUE, ALH and ATHE). Each curve represents a different domain size, with $k$ ranging from $25$ to $10000$. The figure highlights the trade-offs between privacy and adversarial resilience for each protocol, showing how ASR evolves as the privacy budget and domain size change.
@@
General trend across protocols: For all protocols, we observe that the ASR generally increases with increasing privacy budget $\varepsilon$ in Figure 2. This behavior is expected, as higher values of $\varepsilon$ correspond to weaker privacy guarantees, allowing the adversary to infer user data more effectively. For smaller domain sizes (i.e., $k\leq 100$, the ASR rises sharply, suggesting that the adversary’s ability to correctly guess the user’s input improves significantly as $\varepsilon$ increases. In contrast, larger domain sizes (i.e., $k\geq 1000$) show a more gradual increase in ASR, indicating that a larger domain inherently offers greater privacy protection. Nevertheless, our adaptive methods effectively counteract privacy threats regardless of $k$ and $\varepsilon$, underscoring the flexibility and robustness offered by our double-objective optimization framework.
Comparing traditional protocols: Among traditional protocols, GRR and SS demonstrate the highest ASR across all privacy budgets and domain sizes, making them the most vulnerable to privacy attacks. In contrast, SUE displays a more gradual increase in ASR, suggesting better resilience than GRR and SS. Protocols such as OUE and OLH exhibit a maximum ASR of approximately $0.5$, even as $\varepsilon$ increases, showing a clear boundary in their ASR. Notably, for SHE, ASR grows gradually as $\varepsilon$ increases, while THE’s thresholding strategy yields moderate ASR increments. Overall, BLH consistently achieves an exceptionally low ASR across all privacy budgets and domain sizes, highlighting its robustness against adversarial inference attacks.
@@
Our refined and adaptive protocols: Our adaptive protocols (i.e., ASS, AUE, ALH and ATHE) demonstrate significantly lower ASR compared to their traditional counterparts across all privacy budgets. For example, ASS effectively mitigates the ASR vulnerability of the state-of-the-art SS by capping the $\mathrm{ASR}$ below 0.25, in contrast to the traditional SS where the $\mathrm{ASR}$ gets to $1$ (i.e., the adversary can fully infer the user’s value). Similarly, AUE consistently achieves a lower or comparable ASR relative to the state-of-the-art OUE protocol, highlighting the benefits of its adaptive parameter optimization. For ALH, the adaptive mechanism achieves a balanced compromise between the low ASR of BLH and the moderate ASR of OLH by dynamically optimizing the hash domain size, resulting in a substantial reduction in ASR compared to traditional OLH. Finally, ATHE demonstrates more resilience than THE across all domain sizes, thereby showcasing the effectiveness of our framework in enhancing privacy protection.
@@
ASR increases with $\varepsilon$, but our adaptive protocols resist: As the privacy budget $\varepsilon$ increases, ASR generally rises for all protocols due to weaker privacy guarantees. However, our adaptive protocols (ASS, AUE, ALH and ATHE) exhibit significantly lower (i.e., $\leq 5$ orders of magnitude) ASRs across a broad range of $\varepsilon$ values, underscoring their enhanced resilience against privacy attacks and their ability to maintain robust privacy protection.
@@ VI-C MSE Analysis for LDP Protocols
In Figure 3, we examine the MSE behavior of UE-, LH- and HE-based LDP protocols, including our adaptive versions (AUE, ALH and ATHE) across varying privacy budgets $\varepsilon$. In Figure 4, we extend this analysis to compare MSE trends for the SS protocol and its adaptive counterpart ASS. These analyses aim to determine whether the adaptive protocols’ improved robustness to privacy attacks results in substantial estimation accuracy loss or if they maintain competitive MSE values. Notice that the MSE curves of SUE, OUE, BLH, OLH, SHE and THE remain independent of the domain size $k$ due to their fixed parameterization, as established in previous literature [22].
@@
For the SS protocol, we observe in Figure 4 that ASS exhibits an increase in MSE by up to two orders of magnitude compared to SS, particularly in higher privacy regimes ($\varepsilon\geq 4$). This increase is expected, as SS’s minimal MSE comes at the cost of extreme vulnerability, with ASR approaching 1 (e.g., see Figure 2), meaning an adversary can fully infer the user’s value. ASS, in contrast, balances this trade-off by introducing adaptive parameterization that mitigates privacy attacks while moderately increasing the MSE.
In Figure 3, we observe similar trends for the other adaptive protocols (AUE, ALH, and ATHE). For UE protocols, our adaptive AUE version achieves slightly higher MSE compared to OUE across all domain sizes $k$, with the gap becoming more pronounced as $\varepsilon$ increases. This is expected, as AUE optimizes its parameters to enhance robustness to privacy attacks, leading to a slight trade-off in utility. Notably, AUE remains competitive with SUE in high privacy regimes ($\varepsilon\leq 2$), offering similar MSE levels while achieving significantly improved ASR performance as shown earlier. In low privacy regimes ($\varepsilon>2$), AUE incurs a modest increase in MSE, which stays within an order of magnitude compared to OUE.
Moreover, for LH- and HE-based protocols, we observe in Figure 3 that our adaptive protocols (ALH and ATHE) demonstrate a variance (MSE) behavior that is consistently “sandwiched” between the two corresponding state-of-the-art protocols in their respective groups. Specifically, ALH achieves MSE values between those of BLH (which minimizes ASR at the cost of higher MSE) and OLH (which minimizes MSE but is more vulnerable to privacy attacks). Similarly, ATHE’s variance lies between SHE and THE, showing a trade-off where ATHE retains competitive MSE while prioritizing adversarial resilience. This positioning highlights how adaptivity allows our protocols to achieve a better privacy-utility trade-off.
Adaptive protocols maintain competitive MSE: Our adaptive protocols (ASS, AUE, ALH and ATHE) achieve higher MSE compared to their non-adaptive counterparts. However, these increases remain within acceptable bounds ($\leq 2$ orders of magnitude), highlighting that the improved adversarial resilience (ASR $\leq 5$ orders of magnitude) comes at a reasonable cost to the utility. This suggests that enhanced privacy protection against adversaries can be achieved without prohibitive increases in variance.
@@ VI-D Pareto Frontier for ASR and MSE
In addition to evaluating how the ASR and the MSE vary with the privacy budget $\varepsilon$ separately (see Figures 24 and 3), it is insightful to examine how ASR changes as a function of utility loss, measured by the variance (MSE) of the frequency estimation. Figure 5 presents the Pareto frontier between ASR and MSE for the considered LDP protocols, plotting ASR against MSE across small domain sizes and medium to low privacy regimes. This analysis highlights the performance of state-of-the-art protocols (GRR, SUE, BLH, SHE, SS, OUE, OLH and THE) compared to our adaptive variants (ASS, AUE, ALH and ATHE), providing insights into how effectively each method navigates the trade-off between user privacy and data utility. To address space limitations, additional results exploring the ASR-MSE trade-off under varying privacy regimes (high, medium, low) and domain sizes (small, medium, large) are provided in Appendix -C. The findings and discussions in this section are supported by a comprehensive analysis that includes these extended results.
@@
Figure 5: Attacker Success Rate (ASR) vs. Variance (MSE) for numerous LDP frequency estimation protocols. Each plot shows how each protocol performs under varying privacy budgets $\varepsilon$ and domain sizes ($k$), illustrating the trade-off between adversarial success rate (ASR) and utility (MSE). State-of-the-art LDP protocols (i.e., GRR, SUE, BLH, SHE, SS, OUE, OLH and THE) are compared against our adaptive counterparts (i.e., ASS, AUE, ALH and ATHE). Each point represents a different configuration of $\varepsilon$ (in medium to low privacy regimes) and $k$ (small domain), with colors indicating the privacy budget level.
General trend across protocols: For all protocols, there is a clear inverse relationship between ASR and MSE – lower variance estimates correspond to higher ASR, and vice versa. As the privacy budget $\varepsilon$ increases, protocols tend to yield estimates with lower MSE but simultaneously face higher ASRs, reflecting a direct cost to privacy when pursuing higher accuracy. Conversely, under stricter privacy regimes (lower $\varepsilon$), while ASR remains lower, the resulting estimates incur greater MSE. This fundamental tension highlights that simply tuning $\varepsilon$ does not guarantee a balanced privacy-utility outcome. Thus, understanding the ASR-MSE relationship is crucial for informed protocol selection.
Comparing traditional protocols: Among the traditional protocols, GRR and SS tend to cluster in regions with relatively lower MSE but higher ASR, especially at moderate-to-high $\varepsilon$ values. In contrast, SUE, SHE, and THE provide a more gradual trade-off curve, achieving lower ASR at slightly higher MSE levels, suggesting that these methods preserve more privacy when aiming for moderate accuracy. OUE and OLH occupy intermediate positions: they can reach points of low MSE but at the cost of increasing $\varepsilon$. BLH generally exhibits a low ASR while not bounding the MSE, indicating a low privacy-utility trade-off. Overall, none of the traditional protocols dominate the entire ASR-MSE space, and their effectiveness varies with the chosen privacy budget and domain size.
Our refined and adaptive protocols: Our adaptive protocols achieve more favorable ASR-MSE trade-offs, especially providing low-ASR even in low-privacy regimes (i.e., high $\varepsilon$ values). For instance, ASS, derived from SS, effectively prevents the sharp ASR increases observed in SS’s low-MSE regions by fine-tuning protocol parameters, resulting in points that align closer to a Pareto frontier between ASR and MSE. AUE, adapting from OUE, displays a particularly strong trade-off, often settling into low-ASR regimes without disproportionately large MSE. ALH, building on hashing-based approaches, maintains ASR reductions comparable to BLH or OLH but achieves them with more controlled MSE levels, ensuring that the benefits of hashing-based schemes are not compromised. ATHE consistently achieves configurations that lower ASR without excessively increasing MSE, indicating a more harmonious balance.

### v3

@@ V-C1 Adaptive Subset Selection (ASS)
$\begin{aligned} &\underset{\omega\in\mathbb{Z}}{\mathrm{min}}&&J(\omega)=w_{\mathrm{MSE}}\cdot\Bigg(\frac{(k-1)(k+2\omega e^{\varepsilon}-\omega)}{n\omega\left(-k+\omega+(k-1)e^{\varepsilon}-(\omega-1)e^{\varepsilon}\right)^{2}}\\ &&&\left.+\frac{\left(k-\omega+(\omega-1)e^{\varepsilon}\right)\left(-\omega(k-\omega)-\omega(\omega-1)e^{\varepsilon}\right)}{n\omega\left(-k+\omega+(k-1)e^{\varepsilon}-(\omega-1)e^{\varepsilon}\right)^{2}}\right)\\ &&&+w_{\text{ASR}}\cdot\left(\frac{e^{\varepsilon}}{\omega e^{\varepsilon}+k-\omega}\right),\\ &\emph{s.t.}&&1\leq\omega<k,\quad w_{\text{ASR}}+w_{\mathrm{MSE}}=1\mathrm{.}\end{aligned}$
@@
V-C2 Adaptive Unary Encoding (AUE)
@@
$+w_{\text{ASR}}\cdot\frac{1}{k}\left[(1-p)(1-q)^{k-1}+\frac{p}{q}\left(1-(1-q)^{k}\right)\right]\mathrm{,}$
[…]
@@ V-C3 Adaptive LH (ALH)
$\quad+w_{\text{ASR}}\cdot\left(\frac{e^{\varepsilon}}{(e^{\varepsilon}+g-1)\cdot\max\left\{\frac{k}{g},1\right\}}\right)\mathrm{.}$
@@
V-C4 Adaptive THE (ATHE)
The ATHE mechanism extends the THE protocol. Unlike THE [23], which only aims to minimize the MSE in Eq. (24), ATHE jointly optimizes both MSE and ASR by adapting the threshold parameter $\theta$. Formally, the optimization problem for ATHE is defined as:
@@ V-C4 Adaptive THE (ATHE)
[…]
in which $p$ and $q$ are given in Eq. (22). The constraint $0.5\leq\theta\leq 1$ follows the settings used in prior work [23].
V-D Empirical Validation of Analytical ASR-MSE Frontiers
To validate the analytical ASR-MSE framework introduced in Section V-B, we empirically verify that the closed-form ASR and MSE expressions produce Pareto frontiers consistent with the observed behavior on real data. Concretely, we use the Adult dataset from the UCI machine learning repository [48], selecting the Age attribute with domain size $k=100$ (i.e., $\text{Age}\in[0,99]$) and $n=48842$ users, and report results averaged over $100$ independent runs. This experiment serves as a consistency check of the analytical methodology.
[…]
Figure 1 presents the comparison between empirical ($\circ$ markers) and analytical (red dashed lines) Pareto frontiers for ASR versus MSE across various LDP protocols, including both state-of-the-art and our adaptive variants. Each subplot represents a specific protocol (i.e., GRR, SUE, BLH, SHE, SS, OUE, OLH, THE, ASS, AUE, ALH and ATHE) with the privacy budget varying as $\varepsilon\in(2,10)$ with the domain size fixed at $k=100$. One can observe that across all protocols, the empirical results closely align with the analytical calculations, demonstrating the robustness of the closed-form ASR and MSE expressions. This alignment highlights the reliability of our theoretical framework and previous analytical experiments across different privacy budgets and data distributions.
Analytical vs. empirical validation: Our findings demonstrate a strong agreement between analytical and empirical ASR-MSE Pareto frontiers, supporting the use of closed-form expressions as a reliable basis for both systematic analysis and attack-aware LDP protocol design.
@@ VI Experiments and Analysis
The objective of our experiments is to thoroughly evaluate the performance of the proposed adaptive LDP protocols in terms of privacy, utility and resilience against adversarial threats. Specifically, in Section VI-B, we conduct an ASR analysis for each LDP protocol to quantify their vulnerability to data reconstruction attacks, thereby assessing the privacy guarantees offered by existing and newly proposed mechanisms. Subsequently, in Section VI-C, we perform an MSE analysis to evaluate the utility guarantees provided by existing and our newly proposed mechanisms. Next, in Section VI-D, we explore the trade-off between ASR and MSE (i.e., Pareto frontier), highlighting how our adaptive protocols compare to traditional ones in balancing privacy and utility under different scenarios. Finally, in Section VI-E, we extend our evaluation to a joint view of privacy leakage, estimation utility and integrity under data poisoning attacks.
To systematize these evaluations, we rely on analytical closed-form expressions, enabling a comprehensive exploration of the privacy-utility-attackability trade-off across privacy budgets and domain sizes at low computational cost. This choice is supported by prior work showing close agreement between analytical and empirical results [23, 30, 13], and further validated by our empirical study in Section V-D.
@@ VI-A Setup of Analytical Experiments
  • LDP protocols. We experiment with the eight LDP protocols described in Section IV and our four adaptive LDP protocols described in Section V-C.
  • Number of users. For the analytical variance/MSE derived from Eq. (4) (e.g., Eq. (10) for GRR, Eq. (24) for THE, …), we report variance per user, corresponding to the analytical derivation of $\text{Var}[]/n$. This allows us to examine the fundamental properties of each protocol independently of the dataset size.
@@
      High privacy regime with $\varepsilon\in\{0.1,0.2,\ldots,1.9,2.0\}$.
@@
  • Weights for optimization. Unless otherwise mentioned, we fix the weights for the two-objective optimization in Eq. (27) to $w_{\text{ASR}}=w_{\mathrm{MSE}}=0.5$, aiming for a balanced trade-off between privacy and accuracy. Experiments in Appendix -C focus on varying these weights.
  • Optimization method. Parameters for our adaptive protocols (e.g., subset size $\omega$ for ASS) were tuned via grid search [46], followed by selection using a Weighted Scalarization strategy (see Eq. (27)). While grid search is exhaustive and ensures broad coverage of the parameter space, it can be computationally expensive. For large or high-dimensional search spaces, we recommend constrained or gradient-free optimization methods [49] as a more efficient alternative to accelerate convergence without significantly sacrificing solution quality. Experiments in Appendix -D analyze parameter selection.
@@ VI-B ASR Analysis for LDP Protocols
Figure 2: Attacker Success Rate (ASR) vs. privacy budget ($\varepsilon$) for different LDP frequency estimation protocols across varying domain sizes ($k$). The plots compare state-of-the-art LDP protocols, including GRR, SUE, BLH, OUE, OLH, SS, SHE and THE, against our newly proposed adaptive protocols (ASS, AUE, ALH and ATHE). Each curve represents a different domain size, with $k$ ranging from $25$ to $10000$.
@@
For all protocols, we observe that the ASR generally increases with increasing privacy budget $\varepsilon$ in Figure 2. This behavior is expected, as higher values of $\varepsilon$ correspond to weaker privacy guarantees, allowing the adversary to infer user data more effectively. For smaller domain sizes (i.e., $k\leq 100$), the ASR rises sharply, suggesting that the adversary’s ability to correctly guess the user’s input improves significantly as $\varepsilon$ increases. In contrast, larger domain sizes (i.e., $k\geq 1000$) show a more gradual increase in ASR, indicating that a larger domain inherently offers greater privacy protection. Still, our adaptive methods effectively counteract privacy threats regardless of $k$ and $\varepsilon$, underscoring the flexibility and robustness offered by our two-objective optimization framework.
@@
Our adaptive protocols (i.e., ASS, AUE, ALH and ATHE) demonstrate significantly lower ASR compared to their traditional counterparts across all privacy budgets. For example, ASS effectively mitigates the ASR vulnerability of the state-of-the-art SS by capping the $\mathrm{ASR}$ below 0.25, in contrast to the traditional SS, where the $\mathrm{ASR}$ gets to $1$ (i.e., the adversary can fully infer the user’s value). Similarly, AUE consistently achieves a lower or comparable ASR relative to the state-of-the-art OUE protocol, highlighting the benefits of its adaptive parameter optimization. For ALH, the adaptive mechanism achieves a balanced compromise between the low ASR of BLH and the moderate ASR of OLH by dynamically optimizing the hash domain size, resulting in a substantial reduction in ASR compared to traditional OLH. Finally, ATHE demonstrates more resilience than THE across all domain sizes, thereby showcasing the effectiveness of our framework in enhancing privacy protection.
@@
ASR increases with $\varepsilon$, but our adaptive protocols resist: As the privacy budget $\varepsilon$ increases, ASR generally rises for all protocols due to weaker privacy guarantees. However, our adaptive protocols (ASS, AUE, ALH and ATHE) exhibit significantly lower (i.e., by up to $\leq 5\times$) ASRs across a broad range of $\varepsilon$ values, underscoring their enhanced resilience against privacy attacks.
@@ VI-C MSE Analysis for LDP Protocols
In Figure 3, we examine the MSE behavior of the SS protocol and our adaptive protocol ASS, across varying privacy budgets $\varepsilon$. In Figure 4, we extend this analysis to compare MSE trends for UE-, LH- and HE-based LDP protocols, including our adaptive versions (AUE, ALH and ATHE). These analyses aim to determine whether the adaptive protocols’ improved robustness to privacy attacks results in substantial estimation accuracy loss or if they maintain competitive MSE values. Notice that the MSE curves of SUE, OUE, BLH, OLH, SHE and THE remain independent of the domain size $k$ due to their fixed parameterization, as established in [23].
@@
We observe in Figure 3 that ASS exhibits an increase in MSE by up to two orders of magnitude compared to SS, particularly in higher privacy regimes ($\varepsilon\geq 4$). This increase is expected, as SS’s minimal MSE comes at the cost of extreme vulnerability, with ASR approaching 1 (e.g., see Figure 2), meaning an adversary can fully infer the user’s value. ASS, in contrast, balances this trade-off by introducing adaptive parameterization that mitigates privacy attacks while moderately increasing the MSE. In Figure 4, we observe similar trends for the other adaptive protocols (AUE, ALH, and ATHE). For UE protocols, our adaptive AUE version achieves slightly higher MSE compared to OUE across all domain sizes $k$, with the gap becoming more pronounced as $\varepsilon$ increases. This is expected, as AUE optimizes its parameters to enhance robustness to privacy attacks, leading to a slight trade-off in utility. Notably, AUE remains competitive with SUE in high privacy regimes ($\varepsilon\leq 2$), offering similar MSE levels while achieving significantly improved ASR performance as shown earlier. In low privacy regimes ($\varepsilon>2$), AUE incurs a modest increase in MSE (1 order of magnitude) compared to OUE.
Moreover, for LH- and HE-based protocols, we observe in Figure 4 that our adaptive protocols (ALH and ATHE) demonstrate a variance (MSE) behavior that is consistently “sandwiched” between the two corresponding state-of-the-art protocols in their respective groups. Specifically, ALH achieves MSE values between BLH (which minimizes ASR at the cost of higher MSE) and OLH (which minimizes MSE but is more vulnerable to privacy attacks). Similarly, ATHE’s variance lies between SHE and THE, showing a trade-off where ATHE retains competitive MSE while prioritizing adversarial resilience. This positioning highlights how adaptivity allows our protocols to achieve a better privacy-utility trade-off.
Adaptive protocols maintain competitive MSE: Our adaptive protocols (ASS, AUE, ALH and ATHE) achieve higher MSE compared to their non-adaptive counterparts. However, these increases remain within acceptable bounds ($\leq 2$ orders of magnitude), highlighting that the improved adversarial resilience (ASR $\leq 5\times$) comes at a reasonable cost to the utility. This suggests that enhanced privacy protection against adversaries can be achieved without prohibitive increases in variance.
@@ VI-D Pareto Frontier for ASR and MSE
Beyond analyzing ASR and MSE independently as functions of the privacy budget $\varepsilon$ (see Figures 24 and 3), it is crucial to examine how privacy leakage and estimation accuracy interact jointly. To this end, we study the Pareto frontier between attackability (ASR) and estimation error (MSE). Figure 5 reports ASR vs. MSE for a fixed domain size $k=100$, with the privacy budget $\varepsilon$ varying in the medium-to-low privacy regime. Protocols are grouped by their encoding family (GRR/SS, UE, LH and HE), allowing for direct comparison between state-of-the-art mechanisms and their adaptive counterparts under identical domain and privacy conditions. Each point corresponds to a specific value of $\varepsilon$, with colors indicating the privacy level. Additional per-protocol ASR-MSE trade-offs for other domain sizes and privacy regimes are provided in Appendix -E.
@@
Figure 5: Attacker Success Rate (ASR) vs. Variance (MSE) for LDP frequency estimation protocols under data reconstruction attacks, with fixed domain size $k=100$. Each point corresponds to a distinct privacy budget $\varepsilon$ (medium to low privacy regimes), with colors indicating the privacy level. Protocols are organized by encoding family, comparing state-of-the-art mechanisms (GRR, SUE, BLH, SHE, SS, OUE, OLH, THE) against our adaptive counterparts (ASS, AUE, ALH, ATHE).
From Figure 5, comparisons within each encoding family reveal systematic structural differences in how protocols navigate the ASR-MSE trade-off. GRR and SS exhibit steep trajectories: as $\varepsilon$ increases, they rapidly reach low-MSE regimes at the cost of sharply increasing ASR, indicating strong utility but severe privacy leakage. In contrast, SUE-, SHE- and THE-based mechanisms follow smoother ASR-MSE paths, yielding more gradual trade-offs and avoiding abrupt transitions into highly vulnerable regions. OUE and OLH achieve low variance primarily at higher $\varepsilon$ values, where ASR plateaus around $0.5$, reflecting their utility-oriented parameterization. By contrast, BLH consistently maintains low ASR across the entire range of $\varepsilon$, but does so at the expense of substantially higher MSE, illustrating a conservative privacy-first design. Notably, because OUE is designed to minimize MSE, SUE achieves a more favorable ASR-MSE trade-off across a broad range of privacy budgets. This illustrates that minimizing estimation error alone does not necessarily yield the best privacy-utility balance under reconstruction attacks.
Our adaptive protocols fundamentally reshape these within-family trade-offs. ASS mitigates the sharp ASR escalation observed in SS at low-MSE operating points. AUE improves upon OUE by maintaining a lower ASR without incurring disproportionate variance increases, particularly in moderate privacy regimes, and in several regimes, it overlaps with SUE in the ASR-MSE space. Likewise, ALH and ATHE consistently dominate their non-adaptive counterparts by jointly reducing ASR without incurring excessive variance increases. Rather than interpolating between existing designs, these adaptive mechanisms shift the ASR-MSE trade-off frontier itself, enabling operating points that are unattainable through varying $\varepsilon$ alone under fixed protocol parameterizations.